OpenStudy (anonymous):

integrate 1/x^2 -2x + 2 dx

OpenStudy (anonymous):

what to do in the bottom?

OpenStudy (anonymous):

completing the square?

OpenStudy (anonymous):

I would start by completing the square to see if it is an arctan form before doing something like partial fractions. \[\frac{1}{x^2-2x+2}=\frac{1}{x^2-2x+1-1+2}=\frac{1}{(x-1)^2+1}\] \[\int\limits \frac{dx}{(x-1)^2+1}=\arctan(x-1)+C\]

OpenStudy (anonymous):

oh cool.. i was not sure... :)

OpenStudy (anonymous):

No problem :P